OpenStudy (anonymous):

can you tell me which of the following numbers is not a real number -/7 -7 /7 /-7

OpenStudy (phi):

those slash marks mean square root? if so, what is \[ \sqrt{-7} \]

OpenStudy (anonymous):

yeah

OpenStudy (phi):

do you know what is \[ \sqrt{9}= \]? what about \[ \sqrt{-9}= \]?